# CALT-TH 2023-050

### v1

@@ VI.3 Bounds for ${\sf N}=4$
As mentioned in §II, the QEC for ${\sf N}=4$ is not known. However, we will show that our construction, even with a relatively small value of ${\sf N}^{\prime}$, suffices to derive an inner bound that coincides with the subset of $\overline{\Sigma}^{*}_{4}$ of entropy vectors realized by stabilizer states. This subset, which we denote be $\overline{\Lambda}_{4}^{*}$, was shown in Linden et al. (2013) to coincide with its outer bound $\Lambda_{4}$, defined as the polyhedral cone specified by SA, SSA, and Ingleton’s inequality Ingleton (1971). Our goal is to show that there is some $i\in\mathbb{N}$ such that
@@
To see this, we can use a result from Bao et al. (2020a), which constructed hypergraph models realizing the entropy vectors that generate the extreme rays of $\Lambda_{4}$. It is then enough to show that each of these models (seven in total) can be obtained from coarse-graining another hypergraph involving ${\sf N}^{\prime}$ parties (for some ${\sf N}^{\prime}\geq 4$), whose corresponding entropy vector is an extreme ray of the ${\sf N}^{\prime}$-party SAC. Four out of these seven cases are trivial, since they correspond to elements of $\mathscr{R}_{4}$. For the remaining three non-trivial cases (which require $i>0$), we construct the hypergraphs shown in Figures 13. As one can verify, they all are all associated to extreme rays of the SAC for their corresponding number of parties. By an appropriate choice of coarse-grainings, we can reduce these hypergraphs to the ones labeled by 4, 6 and 7 in Figure 4 of Bao et al. (2020a) (up to a trivial permutation of the parties). Specifically, these coarse-grainings are
@@ VII Discussion
For the bounds that we have explicitly constructed in this letter, we relied on the hypergraph models of Bao et al. (2020a) and the result of Walter and Witteveen (2021). However, it was shown in Bao et al. (2020b) that for ${\sf N}\geq 5$ the inclusion of the entropy cone of hypergraphs in the cone of stabilizer states is strict. This implies that hypergraph models might not be sufficient to derive the most stringent possible bounds. On the other hand, we cannot rule out the possibility that hypergraph models do suffice to solve the EQMIP for arbitrary ${\sf N}$, in which case our bounds would not converge to the full QEC for any ${\sf N}\geq 4$. To resolve this, one direction is to look for elements in $\mathscr{R}_{{\sf N}}$ that can be realized by stabilizer states but violate the “hypergraph” inequality found in Bao et al. (2020b) at ${\sf N}=5$.
@@
More generally, it seems conceivable that deep structural properties of the QEC might be extractable by exploring the behavior of the sequence of approximations given in Eq. (6). One possibility, already mentioned in §IV, is that the approximation does not improve beyond a certain point. This was indeed the case for ${\sf N}=3$, but as we have shown, in that case the approximation is exact, and perhaps this is not surprising given that $\overline{\Sigma}^{*}_{3}$ is polyhedral. On the other hand, it is possible that for larger ${\sf N}$, $\overline{\Sigma}^{*}_{{\sf N}}$ is not polyhedral, and it cannot be obtained as the conical hull of a finite set of vectors. Indeed, this is already known to be true for the subset of $\overline{\Sigma}^{*}_{{\sf N}}$ corresponding to classical probability distributions Matús (2007). Is polyhedrality related to the behavior of Eq. (6)? Does the sequence converge? If so, what is its limit, and how is it related to $\overline{\Sigma}^{*}_{{\sf N}}$? We leave these very interesting questions for future work.

### v2

@@ VI.3 Bounds for ${\sf N}=4$
As mentioned in §II, the QEC for ${\sf N}=4$ is not known. However, we will show that our construction, even with a relatively small value of ${\sf N}^{\prime}$, suffices to derive an inner bound that coincides with the subset of $\overline{\Sigma}^{*}_{4}$ of entropy vectors realized by stabilizer states. This subset, which we denote by $\overline{\Lambda}_{4}^{*}$, was shown in Linden et al. (2013) to coincide with its outer bound $\Lambda_{4}$, defined as the polyhedral cone specified by SA, SSA, and Ingleton’s inequality Ingleton (1971). Our goal is to show that there is some $i\in\mathbb{N}$ such that
@@
To see this, we can use a result from Bao et al. (2020a), which constructed hypergraph models realizing the entropy vectors that generate the extreme rays of $\Lambda_{4}$. It is then enough to show that each of these models (seven in total) can be obtained from coarse graining another hypergraph involving ${\sf N}^{\prime}$ parties (for some ${\sf N}^{\prime}\geq 4$), whose corresponding entropy vector is an extreme ray of the ${\sf N}^{\prime}$-party SAC. Four out of these seven cases are trivial, since they correspond to elements of $\mathscr{R}_{4}$. For the remaining three non-trivial cases (which require $i>0$), we construct the hypergraphs shown in Figures 13. As one can verify, they are all associated to extreme rays of the SAC for their corresponding number of parties. By an appropriate choice of coarse grainings, we can reduce these hypergraphs to the ones labeled by 4, 6 and 7 in Figure 4 of Bao et al. (2020a) (up to a trivial permutation of the parties). Specifically, these coarse grainings are
@@ VII Discussion
For the bounds that we have explicitly constructed in this paper, we relied on the hypergraph models of Bao et al. (2020a) and the result of Walter and Witteveen (2021). However, it was shown in Bao et al. (2020b) that for ${\sf N}\geq 5$ the inclusion of the entropy cone of hypergraphs in the cone of stabilizer states is strict. This implies that hypergraph models might not be sufficient to derive the most stringent possible bounds. On the other hand, we cannot rule out the possibility that hypergraph models do suffice to solve the EQMIP for arbitrary ${\sf N}$, in which case our bounds would not converge to the full QEC for any ${\sf N}\geq 4$. To resolve this, one direction is to look for elements in $\mathscr{R}_{{\sf N}}$ that can be realized by stabilizer states but violate the “hypergraph” inequality found in Bao et al. (2020b) at ${\sf N}=5$.
@@
More generally, deep structural properties of the QEC might be extractable by exploring the behavior of the sequence of approximations given in Eq. (6). As we already mentioned at the end of §IV, one possibility (for any given ${\sf N}$) is that there exists some $\widehat{i}$ such that the sequence does not change beyond this point (see Eq. (7)). This is the case for ${\sf N}=3$, where we have shown that the approximation is exact, and the sequence already converges to the full $\overline{\Sigma}^{*}_{3}$ at $i=3$ (see Eq. (10)). However, in principle this behavior is possible even if the approximation is not exact. For example, this would be the case for ${\sf N}=4$ if stabilizer states were sufficient to solve the EQMIP. Another possibility is that (at least for some ${\sf N}$) the sequence continues to grow indefinitely and no such $\widehat{i}$ exists. In that case, it would be important to understand if it converges, and how its limit is related to $\overline{\Sigma}^{*}_{{\sf N}}$.
These possibilities are also related to the question about the polyhedrality of the QEC. For ${\sf N}=3$, the exact approximation of $\overline{\Sigma}^{*}_{3}$ at finite $\widehat{i}$ is possible only because $\overline{\Sigma}^{*}_{3}$ is a polyhedral cone, since any $\Delta_{{\sf N}}^{{\sf N}^{\prime}}$ is polyhedral by construction. For ${\sf N}>3$, it is currently not known whether $\overline{\Sigma}^{*}_{{\sf N}}$ is polyhedral. If it is not, as is the case for the subset of $\overline{\Sigma}^{*}_{{\sf N}}$ corresponding to classical probability distributions Matús (2007), the sequence in Eq. (6) might still converge, but only at infinite $i$.
Finally, let us comment on the role of the outer bound $\Sigma_{{\sf N}}$, and in particular SSA. As we mentioned in the introduction, to our knowledge there is currently no strong evidence against the possibility that the QEC also coincides with $\Sigma_{{\sf N}}$ for ${\sf N}>3$. To explore this possibility, a natural way to proceed is to examine the gap between our inner bound and $\Sigma_{{\sf N}}$. In general, an extreme ray of $\Sigma_{{\sf N}}$ is the intersection of facets corresponding to multiple instances of both SA and SSA, and it may not be an extreme ray of the ${\sf N}$-party SAC. Thus, one may wonder if such an extreme ray can be obtained from coarse graining an extreme ray of the SAC involving more parties and realizable (at least approximately) in quantum mechanics. It is intriguing that an example of this is already present at ${\sf N}=3$. In this case, one of the extreme rays of $\Sigma_{3}$, namely the one corresponding to the four-party GHZ state described in §VI.2, does not saturate any instance of SA, and nevertheless, it can be obtained from our construction via the coarse graining in Eq. (9) of the element of $\mathscr{R}_{6}$ shown in Figure 1. This suggests that further explorations of the gap between our inner bound and the outer bound $\Sigma_{{\sf N}}$ might provide powerful insights on the structure of the quantum entropy cone.